High Country, Inc., produces and sells many recreational products. The company has just opened a new plant to produce a folding camp cot that will be marketed throughout the United States. The following cost and revenue data relate to May, the first month of the plant’s operation:

Beginning inventory 0
Units produced 47,000
Units sold 42,000
Selling price per unit $ 85
Selling and administrative expenses:
Variable per unit $ 3
Fixed (per month) $ 555,000
Manufacturing costs:
Direct materials cost per unit $ 17
Direct labor cost per unit $ 9
Variable manufacturing overhead cost per unit $ 3
Fixed manufacturing overhead cost (per month) $ 893,000

Management is anxious to assess the profitability of the new camp cot during the month of May.

Required:

1. Assume that the company uses absorption costing.

a. Determine the unit product cost.

b. Prepare an income statement for May.

2. Assume that the company uses variable costing.

a. Determine the unit product cost.

b. Prepare a contribution format income statement for May.
Answer #1
1a
Direct materials cost 17
Direct labor cost 9
Variable manufacturing overhead 3
Fixed manufacturing overhead 19 =893000/47000
Unit product cost 48
1b
Absorption Costing Income Statement
Sales 3570000 =42000*85
Cost of goods sold 2016000 =42000*48
Gross profit 1554000
Selling and administrative expenses 681000 =555000+(42000*3)
Net operating income 873000
2a
Direct materials cost 17
Direct labor cost 9
Variable manufacturing overhead 3
Unit product cost 29
2b
Variable Costing Income Statement
Sales 3570000 =42000*85
Variable expenses:
Variable cost of goods sold 1218000 =35000*25
Variable selling and administrative expenses 126000 =42000*3
1344000
Contribution margin 2226000
Fixed expenses:
Fixed manufacturing overhead 893000
Fixed selling and administrative expenses 555000
1448000
Net operating income 778000